# HWC-Loco: A Hierarchical Whole-Body Control Approach to Robust Humanoid Locomotion

### v2

@@ 5 Experiment
Experiment Settings. To conduct a comprehensive evaluation, we quantify the control performance of HWC-Loco in simulated (Issac Gym (Makoviychuk et al., 2021)) and realistic environments from the following perspectives: 1) Scalability: How effectively does HWC-Loco navigate across diverse terrains and adapt to different embodiments? 2) Robustness: How well can HWC-Loco stabilize the humanoid robot under varying levels of disturbance? 3) Naturalness: How effectively does HWC-Loco imitate human-like movement norms? 4) Sensitivity: How promptly can the high-level planning policy activate the appropriate action policy?
@@
Within these experiments, our evaluation metric includes a) Success Rate: The proportion of successful navigation across different scenarios. (Cui et al., 2024) b) Goal Tracking performance: The ability to accurately follow velocity commands by maximizing task rewards $r_{\mathfrak{T}}$ detailed in Appendix A.2 (Chen et al., 2024). c) Human-Like behavior: Measured as the mean squared error (MSE) between the robot’s motions and sampled human motions.
@@
Comparison Methods. To evaluate the effectiveness of different components in HWC-Loco, we design a comparison method using an ablation approach as follows: 1) HWC-Loco-l sets $\alpha$ to a lower value, thereby reducing the sensitivity of the high-level policy to failure events. 2) Goal-Tracking Policy removes the safety recovery mechanism and relies solely on the goal-tracking policy $\pi_{1}$ (Section 4.1). 3) DreamWaQ-Humanoid further removes the human imitation objective $\mathcal{D}_{f}$ from objective (LABEL:obj:goal-tracking), effectively reducing our method to an adaptation of DreamWaQ (Nahrendra et al., 2023) for humanoid control. 4) To highlight the benefits of our hierarchical policy, we compare a recent advancement of DreamWaQ, called AHL (Cui et al., 2024), which employs two-phase training for updating the locomotion policy.
5.1 Scalability: Locomotion across Diverse Terrains
We evaluate humanoid control policies across diverse terrains, including flat ground, obstacles, slopes, and stairs, ensuring comprehensive testing. To enable detailed comparisons, evaluations are conducted under both low-speed and high-speed commands. Specifically, low-speed commands are sampled as $v_{\text{x}}^{c}\sim[0.0,1.0]$, $v_{\text{y}}^{c}\sim[-0.3,0.3]$, and $w_{\text{z}}\sim[-0.5,0.5]$, while high-speed commands follow $v_{\text{x}}^{c}\sim[1.0,2.0]$, $v_{\text{y}}^{c}\sim[-0.6,0.6]$, and $w_{\text{z}}^{c}\sim[-1.0,1.0]$. Here, $v_{\text{x}}^{c}$, $v_{\text{y}}^{c}$, and $w_{\text{z}}$ represent the commanded linear velocities along the x- and y-axes and angular velocity around the z-axis, respectively.
@@ 5.1 Scalability: Locomotion across Diverse Terrains
Table 2 shows that our HWC-Loco achieves the highest success rate across all types of terrain while maintaining promising goal tracking and human mimic performance, demonstrating its exceptional robustness and reliability in the locomotion tasks. Specifically, as the complexity of the terrain increases (from flat surfaces to obstacles, slopes, and stairs), the advantage of our HWC-Loco method in terms of success rate becomes increasingly apparent. Notably, in the high-speed stair terrain setting, HWC-Loco has a significant higher success rate than all other policies, although its goal-tracking ability slightly decreases. This indicates that the policy prioritizes stability over aggressive goal pursuit. For example, when given a high-speed command, HWC-Loco enables the policy to adapt dynamically in safety-critical situations (e.g., navigating stairs mid-air) rather than rigidly maintaining high velocity. Comparably, when downplaying the sensitivity to safe-critical events and removing the safety-recovery policy, the success data drop significantly from nearly 90% to around 60% on the testing environment with stairs and high-speed commands.
Cross Embodiment Validation. We apply HWC-Loco to the Unitree G1 humanoid robot to validate its generalization capability. The testing terrains include all four previously mentioned types, and commands are sampled across the full operational range. The results in Table 1 demonstrate differences between the G1 and H1 across three key metrics.
@@
Unitree H1 $97.13\pm 0.43$ $1.10\pm 0.00$ $1.42\pm 0.01$
Unitree G1 $\boldsymbol{98.14\pm 0.35}$ $\boldsymbol{1.12\pm 0.01}$ $\boldsymbol{1.02\pm 0.02}$
Table 2: Model performance in simulated environments. All policies are trained using three random seeds and then evaluated in 1000 distinct environments. Each evaluation runs for 1200 steps, which is equivalent to 12 seconds of real clock time.
Method Flat Obstacles
(a) Low Speed
DreamWaQ $98.92\pm 0.16$ $1.26\pm 0.01$ $1.50\pm 0.04$ $94.41\pm 0.18$ $1.25\pm 0.02$ $1.50\pm 0.04$
AHL $99.88\pm 0.10$ $1.31\pm 0.02$ $1.42\pm 0.03$ $98.82\pm 0.14$ $1.30\pm 0.01$ $1.42\pm 0.03$
Goal-tracking Policy $99.97\pm 0.01$ $\boldsymbol{1.33\pm 0.01}$ $\boldsymbol{1.39\pm 0.02}$ $99.60\pm 0.10$ $\boldsymbol{1.32\pm 0.01}$ $\boldsymbol{1.39\pm 0.02}$
HWC-Loco-l $\boldsymbol{100.00\pm 0.00}$ $1.32\pm 0.01$ $1.40\pm 0.02$ $\boldsymbol{100.00\pm 0.00}$ $1.32\pm 0.01$ $1.39\pm 0.02$
HWC-Loco $\boldsymbol{100.00\pm 0.00}$ $1.28\pm 0.01$ $1.41\pm 0.03$ $\boldsymbol{100.00\pm 0.00}$ $1.28\pm 0.01$ $1.40\pm 0.02$
(b) High Speed
DreamWaQ $94.38\pm 0.21$ $1.09\pm 0.01$ $1.53\pm 0.08$ $92.21\pm 0.38$ $1.08\pm 0.01$ $1.54\pm 0.09$
AHL $99.78\pm 0.09$ $1.13\pm 0.03$ $1.43\pm 0.03$ $98.38\pm 0.11$ $1.12\pm 0.01$ $1.43\pm 0.04$
Goal-tracking Policy $99.95\pm 0.04$ $\boldsymbol{1.14\pm 0.00}$ $\boldsymbol{1.40\pm 0.03}$ $99.20\pm 0.15$ $\boldsymbol{1.13\pm 0.01}$ $1.42\pm 0.02$
HWC-Loco-l $\boldsymbol{100.0\pm 0.00}$ $1.14\pm 0.02$ $1.41\pm 0.02$ $99.72\pm 0.12$ $1.12\pm 0.03$ $\boldsymbol{1.41\pm 0.03}$
HWC-Loco $\boldsymbol{100.0\pm 0.00}$ $1.13\pm 0.02$ $1.41\pm 0.03$ $\boldsymbol{99.92\pm 0.06}$ $1.12\pm 0.02$ $1.42\pm 0.02$
Method Slops Stairs
(a) Low Speed
DreamWaQ $92.31\pm 0.40$ $1.19\pm 0.02$ $1.52\pm 0.08$ $74.32\pm 1.30$ $1.19\pm 0.03$ $1.56\pm 0.09$
AHL $98.83\pm 0.15$ $1.21\pm 0.01$ $1.42\pm 0.03$ $93.73\pm 0.70$ $1.21\pm 0.04$ $1.42\pm 0.01$
Goal-tracking Policy $99.90\pm 0.04$ $\boldsymbol{1.31\pm 0.00}$ $\boldsymbol{1.40\pm 0.02}$ $96.60\pm 0.14$ $\boldsymbol{1.31\pm 0.00}$ $\boldsymbol{1.40\pm 0.02}$
HWC-Loco-l $\boldsymbol{100.00\pm 0.00}$ $1.23\pm 0.02$ $1.40\pm 0.03$ $99.80\pm 0.08$ $1.21\pm 0.02$ $1.41\pm 0.01$
HWC-Loco $\boldsymbol{100.00\pm 0.00}$ $1.22\pm 0.02$ $1.41\pm 0.06$ $\boldsymbol{99.98\pm 0.01}$ $1.19\pm 0.03$ $1.41\pm 0.03$
(b) High Speed
DreamWaQ $90.46\pm 0.43$ $1.05\pm 0.01$ $1.57\pm 0.10$ $60.58\pm 0.64$ $1.06\pm 0.01$ $1.58\pm 0.12$
AHL $97.36\pm 0.23$ $1.12\pm 0.01$ $1.45\pm 0.03$ $67.48\pm 0.77$ $1.09\pm 0.01$ $1.46\pm 0.04$
Goal-tracking Policy $98.51\pm 0.45$ $\boldsymbol{1.13\pm 0.00}$ $\boldsymbol{1.41\pm 0.03}$ $72.60\pm 0.97$ $\boldsymbol{1.11\pm 0.00}$ $\boldsymbol{1.43\pm 0.07}$
HWC-Loco-l $99.95\pm 0.02$ $1.12\pm 0.02$ $1.42\pm 0.03$ $78.92\pm 0.45$ $1.10\pm 0.01$ $1.44\pm 0.08$
HWC-Loco $\boldsymbol{100.0\pm 0.00}$ $1.14\pm 0.02$ $1.42\pm 0.04$ $\boldsymbol{88.72\pm 0.42}$ $0.99\pm 0.03$ $1.45\pm 0.03$
Table 3: Stability retention rate under impulse disturbances.
Front $\uparrow$ Back $\uparrow$ Side $\uparrow$ Spinning $\uparrow$
DreamWaQ $60.20\pm 0.62$ $57.25\pm 0.52$ $44.44\pm 0.36$ $54.38\pm 0.41$
AHL $90.01\pm 0.64$ $89.15\pm 0.23$ $70.65\pm 0.74$ $82.43\pm 0.24$
G.-t. Policy $94.97\pm 0.54$ $95.15\pm 0.34$ $77.32\pm 0.64$ $88.99\pm 0.14$
HWC-Loco-l $98.35\pm 0.34$ $97.97\pm 0.33$ $88.84\pm 0.20$ $93.95\pm 0.61$
HWC-Loco $\boldsymbol{99.13\pm 0.14}$ $\boldsymbol{99.40\pm 0.30}$ $\boldsymbol{97.90\pm 0.19}$ $\boldsymbol{96.64\pm 0.52}$
5.2 Robustness: Stable Control under Disturbances
We conduct an extensive study on the robustness of HWC-Loco under both simulated and realistic environments.
Robustness under Simulated Disturbance. We evaluated the ability of different methods to resist external impulse disturbances by continuously maintaining a standing posture. The external disturbances include forward forces, backward forces, lateral forces, and spinning torques. As shown in Table 3, HWC-Loco demonstrates exceptional stability, effectively handling both forward and backward forces, with success rates approaching 100%. In response to lateral forces and spinning torques, HWC-Loco also outperforms other policies. Compared to the goal-tracking only policy, these improvements suggest it requires an additional safety recovery policy to regain stability under disturbances.
Robustness in Realistic Deployment. We deploy the HWC-Loco policy on a real humanoid robot and evaluate its performance under various external force disturbances. Specifically, the robot is subjected to continuous interference, including pushing and pulling its arms, kicking its legs, heavier kicks from both the front and back, and the application of rotational torque. Figure 8 (see Appendix B) illustrates the policy-switching process, along with the corresponding changes in the robot’s roll and pitch angles. When encountering external forces, the robot quickly transitions to a recovery policy, adjusting its posture and gait to restore stability. Notably, the policy does not rely solely on the recovery mechanism; instead, it seamlessly switches between the two policies. This behavior highlights the high-level policy’s ability to balance the trade-off between goal tracking and safety recovery.
5.3 Naturalness: Imitating Human-like Movements
In this experiment, we evaluate the robot’s ability to imitate human behavior by adapting its movements in response to varying velocity commands, ranging from low to high values. Figure 7 (see Appendix B) presents the results. When the velocity command is set to 0 m/s, the robot remains stationary. At 1.0 m/s, the robot walks forward with a natural stride, keeping its elbows relatively straight while its arms swing in coordination with its gait. At 2.5 m/s, the robot transitions to running, taking longer strides and gradually bending its elbows, closely resembling human running behavior. As the speed increases, the robot’s stride lengthens, and its step frequency adjusts dynamically, closely mimicking typical human locomotion patterns. This behavior contrasts with some existing approaches (Gu et al., 2024c, d; Cui et al., 2024; Zhang et al., 2024b), where the robot maintains a constant step frequency regardless of the velocity command. In those methods, even at 0 m/s, the robot performs in-place stepping (Zhang et al., 2024b; Cui et al., 2024), which not only appears unnatural but also interferes with tasks requiring stable footing, such as manipulation.
5.4 Sensitivity: Activating the Proper Policy
Sensitive Analysis. We track the averaged proportion of activating goal-tracking policy across 4096 training environments with three random seeds, setting $\alpha$ to 0, 20, 50, and 200. Figure 2 shows the results. For $\alpha=0,20,50$, the goal-tracking policy $\pi_{1}$ remains dominant, with the recovery policy $\pi_{2}$ activating only in extreme cases. However, at $\alpha=200$, training becomes unstable due to sparse termination rewards, causing the recovery policy to be overlooked and leading to overly conservative behavior.
Figure 2: Averaged proportion of goal-tracking policy.
Parkour Test. Figure 9 shows the robot navigates multi-terrain (15cm steps, 20° slopes, wooden boards) at 0.3 m/s without external sensing. Upon encountering challenging terrain, it prioritizes safety by activating a recovery policy instead of persisting with goal tracking. Once it stabilizes on the terrain, it switches back to the goal-tracking policy, demonstrating adaptive control for dynamic environments.

### v3

@@ 5 Experiment
Experiment Settings. To conduct a comprehensive evaluation, we quantify the control performance of HWC-Loco in simulated (Issac Gym [63]) and realistic environments from the following perspectives: 1) Effectiveness: How effectively does HWC-Loco navigate across diverse terrains? 2) Robustness: How well can HWC-Loco stabilize the humanoid robot under varying levels of disturbance? 3) Naturalness: How well does HWC-Loco imitate human-like movement norms? 4) Scalability: How effectively does HWC-Loco generalize to diverse embodiments and tasks?
@@
Our evaluation metric includes a) Success Rate: The proportion of successful traverse across different scenarios. [64] b) Goal Tracking performance: The ability to accurately follow velocity commands by maximizing task rewards $r_{\mathfrak{T}}$ detailed in Appendix A.2 [43]. c) Human-Like behavior: Measured as the Wasserstein-1 distance between the robot’s and sampled human motions.
@@
Comparison Methods. To evaluate the effectiveness of different components in HWC-Loco, we design a comparison method using an ablation approach as follows: 1) HWC-Loco-l sets $\alpha$ to a lower value, thereby reducing the sensitivity of the high-level policy to failure events. 2) Goal-Tracking Policy removes the safety recovery mechanism and relies solely on the goal-tracking policy $\pi_{1}$ (Section 4.1). 3) DreamWaQ-Humanoid further removes the human imitation objective $\mathcal{D}_{f}$ from objective (5), effectively reducing our method to an adaptation of DreamWaQ [18] for humanoid control. 4) To highlight the benefits of our hierarchical policy, we compare a recent advancement of DreamWaQ, called AHL [64], which employs two-phase training for updating the locomotion policy. All policies are trained using three random seeds and evaluated in 1000 distinct environments.
5.1 Effectiveness: Locomotion across Diverse Terrains
@@
Effectiveness in Simulated Environments. We evaluate the policies on challenging terrains, specifically slopes and stairs. To enable a comprehensive comparison, evaluations are performed under both low-speed and high-speed command conditions. Low-speed commands are sampled as $v_{\text{x}}^{c}\sim[0.0,1.0]$, $v_{\text{y}}^{c}\sim[-0.3,0.3]$, and $w_{\text{z}}^{c}\sim[-0.5,0.5]$, while high-speed commands follow $v_{\text{x}}^{c}\sim[1.0,2.0]$, $v_{\text{y}}^{c}\sim[-0.6,0.6]$, and $w_{\text{z}}^{c}\sim[-1.0,1.0]$. Here, $v_{\text{x}}^{c}$ and $v_{\text{y}}^{c}$ denote the commanded linear velocities along the x- and y-axes, respectively, and $w_{\text{z}}^{c}$ denotes the commanded angular velocity around the z-axis. Table 1 shows that our HWC-Loco achieves the highest success rate across all evaluated scenarios, while maintaining promising goal tracking and human mimic performance. Notably, in the high-speed stair terrain setting, HWC-Loco has a significantly higher success rate than all other policies, although its goal-tracking ability slightly decreases. This indicates that the policy prioritizes stability over aggressive goal pursuit. For example, when given a high-speed command, HWC-Loco enables the policy to adapt dynamically in safety-critical situations (e.g., navigating stairs mid-air) rather than rigidly maintaining high velocity. Comparably, when downplaying the sensitivity to safety-critical events and removing the safety-recovery policy, the success data drops significantly from nearly 85% to around 60% in the testing environment with stairs and high-speed commands.
Effectiveness in Realistic Environments. Figure 10 (see Appendix B.1) demonstrates the robot’s ability to climb 15 cm stairs and 20° slopes. When faced with challenging terrain, the robot prioritizes safety by activating the recovery policy. After regaining stability, it smoothly transitions back to goal tracking, showcasing its adaptive control in dynamic and uncertain environments. Figure 11 (see Appendix B.1) further illustrates the robot’s effectiveness in real-world outdoor settings, where it successfully navigates across diverse terrains, including flat ground, grassy surfaces, and slopes.
Table 1: Locomotion performance in simulated environments. Each evaluation runs for 1200 steps, which is equivalent to 12 seconds of real clock time.
Method Slopes Stairs
(a) Low Speed
DreamWaQ $92.31\pm 0.40$ $1.19\pm 0.02$ $3.52\pm 0.05$ $74.32\pm 1.30$ $1.19\pm 0.03$ $3.53\pm 0.02$
AHL $98.83\pm 0.15$ $1.21\pm 0.01$ $3.41\pm 0.04$ $93.73\pm 0.70$ $1.21\pm 0.04$ $3.44\pm 0.01$
Goal-tracking Policy $99.90\pm 0.04$ $\boldsymbol{1.31\pm 0.00}$ $\boldsymbol{3.30\pm 0.05}$ $96.60\pm 0.14$ $\boldsymbol{1.31\pm 0.00}$ $\boldsymbol{3.31\pm 0.01}$
HWC-Loco-l $\boldsymbol{100.00\pm 0.00}$ $1.23\pm 0.02$ $3.32\pm 0.03$ $99.80\pm 0.08$ $1.21\pm 0.02$ $3.31\pm 0.06$
HWC-Loco $\boldsymbol{100.00\pm 0.00}$ $1.22\pm 0.02$ $3.32\pm 0.05$ $\boldsymbol{99.98\pm 0.01}$ $1.19\pm 0.03$ $3.34\pm 0.03$
(b) High Speed
DreamWaQ $90.46\pm 0.43$ $1.05\pm 0.01$ $3.63\pm 0.04$ $60.58\pm 0.64$ $1.06\pm 0.01$ $3.66\pm 0.02$
AHL $97.36\pm 0.23$ $1.12\pm 0.01$ $3.76\pm 0.02$ $67.48\pm 0.77$ $1.09\pm 0.01$ $3.68\pm 0.04$
Goal-tracking Policy $98.51\pm 0.45$ $\boldsymbol{1.13\pm 0.00}$ $3.52\pm 0.03$ $72.60\pm 0.97$ $\boldsymbol{1.11\pm 0.00}$ $3.43\pm 0.02$
HWC-Loco-l $99.95\pm 0.02$ $1.12\pm 0.02$ $\boldsymbol{3.43\pm 0.04}$ $78.92\pm 0.45$ $1.10\pm 0.01$ $\boldsymbol{3.39\pm 0.05}$
HWC-Loco $\boldsymbol{100.0\pm 0.00}$ $1.12\pm 0.02$ $3.44\pm 0.09$ $\boldsymbol{84.34\pm 0.43}$ $1.07\pm 0.03$ $3.46\pm 0.03$
5.2 Robustness: Stable Control under Disturbances
[…]
As shown in Table 2, HWC-Loco consistently achieves the highest success rates across all types of disturbances. Specifically, under constant disturbances, HWC-Loco demonstrates superior performance with a success rate of 75.95%, which is nearly 25% higher than that of DreamWaQ. It validates the effectiveness of HWC-Loco’s recovery mechanism in resisting strong and persistent forces. Under low-impulse disturbances, all methods maintain relatively high success rates, with HWC-Loco again achieving the best performance at 94.84%. When facing more challenging high-impulse disturbances, the performance gap becomes more pronounced. HWC-Loco reaches a success rate of 81.27%, outperforming all baselines by a significant margin. This result indicates that the recovery mechanism effectively mitigates the impact of sudden impulses by enabling the robot to quickly lower its body and wave its arms to regain stability. In the payload test, HWC-Loco also demonstrates the best performance. Although it s not explicitly trained on scenarios such as adding 10 kg to the hands, it maintains a high success rate. This highlights the robustness of HWC-Loco to unseen disturbances.
Table 2: Robustness of Locomotion across Various Terrains under Different Disturbances
Policy External Force/Torque Disturbances Impulse Disturbances on CoM Payload on Upper Body
Low-freq. $\uparrow$ Constant $\uparrow$ Low-imp. $\uparrow$ High-imp. $\uparrow$ Low Payload $\uparrow$ High Payload $\uparrow$
DreamWaQ $85.92\pm 1.64$ $51.31\pm 0.97$ $85.24\pm 0.37$ $45.34\pm 0.41$ $67.63\pm 3.10$ $55.04\pm 2.82$
AHL $87.15\pm 2.86$ $60.72\pm 0.36$ $85.87\pm 1.47$ $62.94\pm 1.29$ $79.29\pm 0.65$ $59.16\pm 1.60$
Goal-tracking Policy $90.00\pm 2.20$ $61.20\pm 0.61$ $88.90\pm 0.60$ $58.13\pm 0.78$ $78.34\pm 1.23$ $61.44\pm 1.00$
HWC-Loco-l $92.69\pm 0.56$ $68.60\pm 0.28$ $92.79\pm 1.75$ $77.09\pm 0.79$ $84.11\pm 2.66$ $63.96\pm 0.09$
HWC-Loco $\boldsymbol{95.88\pm 0.37}$ $\boldsymbol{75.95\pm 0.66}$ $\boldsymbol{94.84\pm 0.54}$ $\boldsymbol{81.27\pm 0.80}$ $\boldsymbol{87.43\pm 0.92}$ $\boldsymbol{69.86\pm 1.00}$
Robustness in Realistic Deployment. We deploy HWC-Loco on a real humanoid robot and evaluate its performance under various external force disturbances. Figure 12 (see Appendix B.1) shows policy switching and associated changes in roll and pitch angles under continuous perturbations, including pushes, pulls, and kicks. Upon encountering disturbances, the robot promptly switches to the recovery policy, adjusting its posture and gait to regain stability. Importantly, the controller does not rely solely on recovery mode but dynamically switches between goal-tracking and recovery policies, thereby adapting the action distribution to environmental changes. Figure 13 (see Appendix B.1) further demonstrates the robot’s robustness to disturbances in outdoor settings.
5.3 Naturalness: Imitating Human-like Movements
This experiment assesses the robot’s ability to imitate human-like locomotion under varying velocity commands. As shown in Figure 4, the robot remains stationary at 0 m/s, and exhibits natural walking behavior at 1.0 m/s, characterized by straight elbows and coordinated arm swings. At 2.5 m/s, it transitions to running, with longer strides and increasingly bent elbows, closely resembling human running. As the velocity increases, both stride length and step frequency adjust dynamically, mimicking typical human locomotion. In contrast, prior approaches [28, 67, 64, 68] often maintain a fixed step frequency regardless of the commanded velocity. Some even exhibit in-place stepping at 0,m/s [68, 64], resulting in unnatural behavior for tasks requiring a steady stance, such as manipulation.
5.4 Scalability: Extending HWC-Loco to New Tasks
Cross Embodiment Validation. We deploy HWC-Loco on the Unitree G1 humanoid robot to evaluate its generalization capability under the same locomotion settings. As shown in Table 3, G1 outperforms H1 across all metrics. A detailed analysis is provided in Appendix A.2.
Figure 4: Human-like behavior.
@@ 5.4 Scalability: Extending HWC-Loco to New Tasks
Unitree H1 $97.13\pm 0.43$ $1.10\pm 0.00$ $3.18\pm 0.01$
Unitree G1 $\boldsymbol{98.14\pm 0.35}$ $\boldsymbol{1.12\pm 0.01}$ $\boldsymbol{3.11\pm 0.03}$
Table 4: Motion Tracking under Disturbances
Punching $\uparrow$ Dancing $\uparrow$ Expressive Walking $\uparrow$
HWC-Loco $\boldsymbol{94.01\pm 0.49}$ $\boldsymbol{86.44\pm 0.60}$ $\boldsymbol{94.53\pm 0.23}$
Motion-tracking $90.85\pm 0.74$ $81.64\pm 0.64$ $89.63\pm 0.27$
Expressive Motion Tracking We follow the approach of ExBody[33, 69] and set expressive motion as the goal for HWC-Loco. We train the model following the same pipeline as locomotion tasks. A domain-randomized motion tracking policy serves as the baseline. We evaluate both policies on three representative motions under impulse disturbances. The success rates are reported in Table 4.
6 Limitation
Our approach has three main limitations. First, policy transitions are governed by a learned discrete decision module, while the low-level controllers remain fixed. A jointly optimized hierarchical framework may enable smoother and more adaptive transitions. Second, the humanoid robot used in real-world deployment has only 19 degrees of freedom, which limits whole-body coordination and constrains the expression of complex recovery behaviors. Third, the recovery policy is trained in simulated dynamic scenarios that may not fully reflect the diversity and complexity of real-world disturbances. Incorporating more realistic perturbations, such as those generated by adversarial policies, could further improve recovery ability.